simplify these expressions
(3+2i)(3-2i)
8/-2i
: simplify these expressions
(3+2i)(3-2i)
8/-2i

Answer by orca(336)

(3+2i)(3-2i)
Solution:
Using the formula (a+b)(a-b)=a^2 - b^2 to expand (3+2i)(3-2i), we have
(3+2i)(3-2i)
=3^2 - (2i)^2
= 9 - 2^2*i^2
= 9 - 4*(-1)
= 9 + 4
= 13
***************************************************************************
8/-2i
Solution:
Our goal is to remove i from the denominator. It can be achieved by multiply both the denominator and numerator by i.
8/(-2i)
= 8*i/(-2i*i)
= 8i/(-2*i^2)
= 8i/[-2*(-1)]
= 8i/2
= 4i